OpenStudy (anonymous):

true or false Three triangles have sides of lengths 3, 4 and 5. Their respective perimeters are 6, 8 and 10. The triangles are similar to each other.

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

6/3 = 8/4 = 10/5 = that should tell you

OpenStudy (mertsj):

3,4,5 is a pythagorean triple and 6,8,10 is a multiple of that. Also the corresponding sides have equal ratios.